A 5.00-g lead bullet moving at 400 m/s penetrates a wood block and comes to rest inside the block. If half of its kinetic energy goes into heating the bullet, what is the bullet's change in temperature? The specific heat of lead is 128 J/ (kg*K).
Answer #1

m=5g=0.005 kg,v=400 m/s

kinetic energy of bullet=0.5*m*v^2=400 J

heating energy=400/2=200 J

by calorimeter principle

heat energy=mass*specific heat*change in temperature

200=0.005*128*change in temperature

change in temperature=312.5 K